OpenStudy (anonymous):

marcy has a total of 100 dimes and quarters. If the total value of the coins is $14.05, how many quarters does she have?

OpenStudy (anonymous):

x = the number of dimes y = the number of quarters x + y = 100 .10x + .25y = 14.05 x + y = 100 10x + 25y = 1405 -10(x + y = 100) 10x + 25y = 1405 -10x - 10y = -1000 10x + 25y = 1405 15y = 405 y = 27 73 dimes 27 quarters

OpenStudy (anonymous):

thank u soo much! (;